# Inequality Restricted Minimum Density Power Divergence Estimation in Panel Count Data

### v1

@@ 5 Optimal tuning Parameter
$\left.+\left[t\left(\frac{p\left(\bm{n}\mid\bm{\theta}\right)}{p\left(\bm{n}^{(s-)}\mid\bm{\theta}\right)}\right)-t\left(\frac{q\left(\bm{n}\right)}{q\left(\bm{n}^{(s-)}\right)}\right)\right]^{2}\right\},$ (5.3)
Based on (5.3), the empirical estimator of the objective function is given by
@@
$\quad-2t\left(\frac{p\left(\bm{n}^{(s+)}\mid\bm{\theta}\right)}{p\left(\bm{n}\mid\bm{\theta}\right)}\right)\Bigg{]}$ (5.4)
The optimal tuning parameter $\gamma$ is then derived as,
${\hat{\bm{\gamma}}}=\underset{\bm{\theta}}{\arg\min}\ \widehat{D}_{\mathrm{GSM}}\left(q,p\right),$ (5.5)
@@ 6.1. Simulation Study
In this section, a simulation environment is developed exploiting 1000 Monte Carlo generations to evaluate the performance of MLE and MDPDEs. Additionally, a comparative analysis is performed to examine the effectiveness of these estimators in both restricted and unrestricted scenarios. The number of occurrences of the recurrent events is considered to follow the Non-homogeneous poison process. The true values of the model parameters are assigned as $\zeta=4.5$, $a_{1}=0.9$, $b_{1}=0.5$, $a_{2}=0.6$, $b_{2}=0.2$, and the time intervals are defined as $(0.01,0.35]$, $(0.35,0.69]$, and $(0.69,1.12]$. The experiment is performed with a fixed sample size $m=100$ and tuning parameters $\gamma=0.2,0.5,0.8$.
Moreover, the behavior of the estimators has been thoroughly examined for both pure and contaminated data. Contamination is incorporated in the data set by generating an $\epsilon$ proportional of values from the Inverse-gaussian distribution. In this context, we examine four distinct contamination rates, $\epsilon=0.153,0.085,0.044,0.006$, and analyze the results accordingly. Here, we have utilized the Sequential convex programming (SCP) algorithm to obtain the MLE and MDPDEs.
In both restricted and unrestricted scenarios, it has been observed that MLE exhibits lower bias and MSE than MDPDEs when dealing with pure data, making it the superior choice in such cases. However, under various levels of data contamination, MDPDEs show reduced bias and MSE in comparison to MLE. Moreover, the increase in bias for MDPDEs from pure to contaminated data is less significant than that for MLE. This indicates that MDPDE is a robust estimator that does not significantly compromise efficiency. The graphical representations of bias estimates for model parameters in both pure and contaminated datasets under restricted and unrestricted scenarios are provided in Figure 1 and Figure 2, respectively. Owing to the superiority of MDPDE on the basis of performance demonstrated here; it is a preferable choice over MLE when the data is contaminated.
@@
Further, the optimal behavior of the tuning parameter $\gamma$ is studied for pure data and contaminated data. In addition to finding the value of $\gamma$, the MSE of the model parameters is calculated under varying contamination levels with a fixed sample size of $m=100$. This same analysis has been conducted for both the GSM and IWJ methods for a comparison purpose. Table 3 and Table 4 show the optimal tuning parameters identified through the GSM and IWJ approaches at $\bm{\epsilon}=0.000,0.006,0.044,0.085,0.153$. According to these optimal values of $\gamma$, the MSE of the model parameters is calculated to assess the performance of each method. The results reveal that the MSE obtained through the GSM method is significantly lower than that achieved using the IWJ method, indicating the superior accuracy and effectiveness of the GSM approach in identifying the optimal tuning parameter. Consequently, it can be inferred that GSM provides a more effective and reliable approach for finding the optimal tuning parameter $\gamma$ compared to the IWJ method.
@@
Table 1: MSE of estimates of parameters under Unrestricted case
                 $\bm{\zeta}$ 0.017465 0.039252 0.076559 0.087280 0.101748
                 a1 0.003021 0.005479 0.007458 0.009799 0.012026
MLE        b1 0.007081 0.010134 0.012235 0.014341 0.016069
                 a2 0.007263 0.009911 0.011577 0.014249 0.016204
                 b2 0.008687 0.011307 0.013132 0.015194 0.017327
                 $\bm{\zeta}$ 0.021666 0.022318 0.022976 0.023651 0.024617
                 a1 0.003563 0.003798 0.004042 0.004308 0.004550
$\bm{\gamma=0.2}$    b1 0.007732 0.008452 0.008908 0.009285 0.009916
                 a2 0.008061 0.008379 0.008720 0.009093 0.009522
                 b2 0.009093 0.009520 0.009951 0.010257 0.010811
                 $\bm{\zeta}$ 0.022637 0.023190 0.023959 0.024920 0.025759
                 a1 0.003347 0.003544 0.003773 0.004042 0.004239
$\bm{\gamma=0.5}$    b1 0.007362 0.007625 0.007941 0.008624 0.009454
                 a2 0.007731 0.007974 0.008413 0.008650 0.009104
                 b2 0.009404 0.009628 0.010179 0.010594 0.011111
                 $\bm{\zeta}$ 0.020219 0.020992 0.021446 0.022011 0.022565
                 a1 0.003050 0.003341 0.003519 0.003728 0.004096
$\bm{\gamma=0.8}$    b1 0.007794 0.008007 0.008599 0.008959 0.009685
                 a2 0.007896 0.008183 0.008531 0.008857 0.009219
                 b2 0.008947 0.009331 0.009605 0.010118 0.010461
@@
                 $\bm{\zeta}$ 0.009932 0.034472 0.053521 0.065142 0.084081
                 a1 0.000466 0.001334 0.001936 0.002625 0.004339
MLE        b1 0.005420 0.008088 0.009605 0.012155 0.015368
                 a2 0.005231 0.008586 0.010277 0.012872 0.014855
                 b2 0.006134 0.007892 0.011879 0.013793 0.016560
                 $\bm{\zeta}$ 0.010005 0.010065 0.010196 0.010352 0.010598
                 a1 0.000506 0.000598 0.000699 0.000797 0.000939
$\bm{\gamma=0.2}$    b1 0.006012 0.006434 0.006883 0.007154 0.007562
                 a2 0.005637 0.006068 0.006330 0.006779 0.007383
                 b2 0.005823 0.006152 0.006688 0.007104 0.007546
                 $\bm{\zeta}$ 0.010430 0.010717 0.010840 0.011036 0.011273
                 a1 0.000554 0.000654 0.000761 0.000842 0.001081
$\bm{\gamma=0.5}$    b1 0.005816 0.006217 0.006548 0.006975 0.007377
                 a2 0.005474 0.005726 0.005933 0.006281 0.006584
                 b2 0.005724 0.005975 0.006444 0.006722 0.007127
                 $\bm{\zeta}$ 0.010707 0.010956 0.011433 0.011871 0.012092
                 a1 0.000591 0.000689 0.000897 0.000984 0.001234
$\bm{\gamma=0.8}$    b1 0.005751 0.006028 0.006285 0.006715 0.006988
                 a2 0.005504 0.005874 0.006086 0.006487 0.007082
                 b2 0.005621 0.005806 0.006182 0.006405 0.006792
@@
              $\bm{\gamma}$ 0.44 0.19 0.41 0.23
              $\bm{\zeta}$ 0.010456 0.010589 0.010401 0.010457
              a1 0.000489 0.000512 0.000537 0.000559
MSE    b1 0.005725 0.005912 0.005642 0.006421
              a2 0.006213 0.006324 0.006143 0.007254
              b2 0.005897 0.006019 0.006163 0.007462
6.2. Real Data Analysis
This research examines survey data from the National Health and Nutrition Examination Survey (NHANES) carried out from 2017 until March 2020. As a result of the COVID-19 pandemic, field operations for the 2019–2020 cycle were suspended, resulting in data that is incomplete and not representative of the nation. To tackle this issue, a nationally representative dataset was formed by merging data from the 2017–2018 cycle with pre-pandemic information from the 2019–March 2020 cycle. This data is retrieved from the adjacent link, National Health and Nutrition Examination Survey (NHANES); NHANES 2017-March 2020 Pre-pandemic Questionnaire Data - https://www.cdc.gov/nchs/nhanes. Here, the study centers on 50 subjects with equivalent follow-up durations, exploring $N_{i1l}$ and $N_{i2l}$, which represent the number of cold/flu and pneumonia events, correspondingly, in the period $(\tau_{l-1},\tau_{l}]$. These subjects have been observed every 3 months, within an entire period of 12 months.
The real dataset reveals a few outliers, particularly in the relationship between frequent cold/flu and the likelihood of developing pneumonia. In most cases, individuals who have experienced cold or flu episodes four to five times within a specific interval tend to develop pneumonia at least once in that period. However, 6 out of 50 individuals deviate from this pattern. Despite suffering from cold/flu five to six times in a given timeframe, they did not develop pneumonia at all. A key observation is that these individuals engage in a higher-than-average duration of physical activity, as depicted in the stacked bar graph (Figure 3), which visualizes the recurrence of cold/flu ($N_{i1l}$) and pneumonia ($N_{i2l}$) across individuals. The frailty variable $Z_{i}$, representing the number of hours a person engages in physical activities weekly, emerges as a crucial determinant in this anomaly. It appears that higher physical activity levels serve as a protective factor, preventing pneumonia during the follow-up period. This correlation is further illustrated in the bubble chart (Figure 4), where these six individuals are represented by large bubbles, reflecting their high exercise rates. Keeping in view the current scenario, the robust restricted minimum density power divergence (RMDPD) may be a suitable estimation technique.
@@ 6.2. Real Data Analysis
Further, a dissimilarity measure-based test statistic to evaluate the suitability of the nonhomogeneous Poisson process for the data is defined as,
$T=\frac{1}{mk}\sum_{i=1}^{m}\sum_{l=1}^{k}\left[\left|\frac{n_{i1l}-e_{i1l}}{e_{i1l}}\right|+\left|\frac{n_{i2l}-e_{i2l}}{e_{i2l}}\right|\right],$ (6.1)
where $n_{i1l}$ and $e_{i1l}$ represent the observed and expected counts of cold or flu occurrences for the $i^{\text{th}}$ individual during the interval $(\tau_{l-1},\tau_{l}]$, respectively, and $n_{i2l}$ and $e_{i2l}$ represent the observed and expected counts of pneumonia occurrences for the same individual within the interval $(\tau_{l-1},\tau_{l}]$, with $i=1,2,3,\ldots,50$. The bootstrap (Bt) technique is employed to estimate the p-value of the test using 10,000 generated samples. A test statistic is computed for each sample. The proportion of bootstrap test statistics that exceed the test statistic from the real data, T, divided by the total number of bootstrap samples, provides the approximate p-value. A smaller test statistic indicates a greater likelihood that the data conform to the nonhomogeneous Poisson process. The test statistic along with the associated p-value for various tuning parameters $(\gamma=0.2,0.5,0.8)$ are reported in Table 5. These p-values confirm the suitability of the nonhomogeneous Poisson process for modeling data. Therefore, it can be inferred that the instances of cold or flu and pneumonia in people can be adequately represented by the nonhomogeneous Poisson process.
@@
0.2 4.804327 0.683200
0.5 4.862886 0.639600
0.8 5.005534 0.544300
The MLE and RMDPDE of the model parameters are obtained, and the bootstrap estimates of Bias and MSE generated from 10,000 bootstrap samples are computed. The initial parameter value of the Sequential convex programming algorithm is set as (6.75, 0.82, 0.63, 0.42, 0.23), which is selected through a comprehensive grid search method. It is noticed that MDPDE shows lower Bootstrap Bias and MSE (Bt Bias and Bt MSE) in comparison to MLE. Considering that the dataset contains some outliers, the efficacy of RMDPDE in this context aligns with the aim of the current study. Consequently, the restricted MDPDE has demonstrated outstanding performance on the real dataset. The Bootstrap Bias and MSE (Bt Bias and Bt MSE) of MLE and RMDPDE with different $\gamma$ values are presented in Table 6.
Table 6: Bootstrap Bias and MSE of estimates of parameters
@@
Bt Bias MLE 0.025789 0.069823 0.054001 -0.005962 -0.005547
$\bm{\gamma=0.2}$ 0.009465 0.033071 0.021024 -0.003395 -0.002130
$\bm{\gamma=0.5}$ 0.009234 0.027456 0.018563 -0.002843 -0.002258
$\bm{\gamma=0.8}$ 0.009021 0.019354 0.011279 -0.001865 -0.002389
Bt MSE MLE 0.000678 0.003456 0.006283 0.000049 0.000031
$\bm{\gamma=0.2}$ 0.000090 0.001094 0.000478 0.000016 0.000012
$\bm{\gamma=0.5}$ 0.000086 0.000876 0.000379 0.000011 0.000014
$\bm{\gamma=0.8}$ 0.000082 0.000475 0.000139 0.000008 0.000015
The optimal tuning parameter, $\gamma$, for the model, is determined using real data. Based on this value of $\gamma$, the model parameters $\zeta$, $a_{1}$, $b_{1}$, $a_{2}$, and $b_{2}$ are estimated, and the Bias and mean squared error (MSE) of the parameters are calculated. The results obtained using GSM method are compared with those from the IWJ method. Table 7 provides a detailed summary of the optimal tuning parameter $\gamma$, the estimated model parameters, and their corresponding Bias and MSE values. It has been noted that the estimates derived from the GSM method consistently show lower Bt Bias and Bt MSE compared to those obtained from the IWJ method. This comparison reveals that the GSM method outperforms the IWJ method by achieving lower MSE, underscoring its effectiveness in tuning parameter selection.

### v4

@@ 5 Optimal tuning Parameter
$\left.+\left[t\left(\frac{p\left(\bm{n}\mid\bm{\theta}\right)}{p\left(\bm{n}^{(s-)}\mid\bm{\theta}\right)}\right)-t\left(\frac{q\left(\bm{n}\right)}{q\left(\bm{n}^{(s-)}\right)}\right)\right]^{2}\right\}.$ (21)
Based on (21), the empirical estimator of the objective function is given by
@@
$\quad-2t\left(\frac{p\left(\bm{n}^{(s+)}\mid\bm{\theta}\right)}{p\left(\bm{n}\mid\bm{\theta}\right)}\right)\Bigg].$ (22)
The optimal tuning parameter $\gamma$ is then derived as
${\hat{\bm{\gamma}}}=\underset{\bm{\theta}}{\arg\min}\ \widehat{D}_{\mathrm{GSM}}\left(q,p\right).$ (23)
@@ 6.1. Simulation Study
In this section, a simulation environment is developed exploiting 1000 Monte Carlo generations to evaluate the performance of MLE and MDPDEs. Additionally, a comparative analysis is performed to examine the effectiveness of these estimators in both restricted and unrestricted scenarios. The number of occurrences of the recurrent events is considered to follow the non-homogeneous Poisson process. In our formulation, we assume that the recurrent event $N_{i1l}$ occurs more frequently than $N_{i2l}$. To reflect this assumption in the model, we impose the inequality constraints $a_{1}\geq a_{2}$ and $b_{1}\geq b_{2}$, ensuring that the rate parameters associated with $N_{i1l}$ are at least as large as those of $N_{i2l}$. These constraints help incorporate the inherent ordering of event frequencies into the estimation process. The true values of the model parameters are assigned as $\zeta=4.5$, $a_{1}=0.9$, $b_{1}=0.5$, $a_{2}=0.6$, $b_{2}=0.2$, and the time intervals are defined as $(0.01,0.35]$, $(0.35,0.69]$, and $(0.69,1.12]$. The experiment is performed with a fixed sample size $m=100$ and tuning parameters $\gamma=0.2,0.5,0.8$.
Moreover, the behavior of the estimators has been thoroughly examined for both pure and contaminated data. Contamination is incorporated in the data set by generating an $\epsilon$ proportional of values from the Inverse-Gaussian distribution. In this context, we examine four distinct contamination rates, $\epsilon=0.006,0.044,0.085,0.153$, and analyze the results accordingly. Here, we have utilized the sequential convex programming (SCP) algorithm to obtain the MLE and MDPDEs.
An examination of the parameter estimates reported in Table 1 and Table 2 reveals that the inequality constraints $a_{1}\geq a_{2}$ and $b_{1}\geq b_{2}$ are consistently satisfied under both the unrestricted and restricted estimation schemes.
Further, in both unrestricted and restricted scenarios, it has been observed that MLE exhibits lower bias and MSE than MDPDEs when dealing with pure data, making it the superior choice in such cases. However, under various levels of data contamination, MDPDEs show reduced bias and MSE in comparison to MLE. Moreover, the increase in bias for MDPDEs from pure to contaminated data is less significant than that for MLE. This indicates that MDPDE is a robust estimator that does not significantly compromise efficiency. The graphical representations of bias estimates for model parameters in both pure and contaminated datasets under restricted and unrestricted scenarios are provided in Figure 1 and Figure 2, respectively. Owing to the superiority of MDPDE on the basis of performance demonstrated here, it is a preferable choice over MLE when the data is contaminated.
@@
Table 1: Estimated values of the parameters under Unrestricted case
        $\bm{\zeta}$ 4.599616 4.685667 4.731346 4.755229 4.789967
        a1 0.954962 0.974019 0.986361 0.998989 1.009662
MLE     b1 0.584146 0.600669 0.610611 0.619753 0.626762
        a2 0.684146 0.700669 0.710611 0.719753 0.726762
        b2 0.293205 0.306336 0.314596 0.323264 0.331632
        $\bm{\zeta}$ 4.600026 4.600325 4.600974 4.601743 4.602945
        a1 0.959689 0.960625 0.961580 0.962635 0.963452
$\bm{\gamma=0.2}$    b1 0.587931 0.589936 0.591382 0.593357 0.595581
        a2 0.686931 0.687936 0.689382 0.690357 0.692581
        b2 0.295359 0.296568 0.298754 0.300279 0.301976
        $\bm{\zeta}$ 4.602129 4.603525 4.604113 4.605053 4.606175
        a1 0.957855 0.959529 0.960423 0.963579 0.964110
$\bm{\gamma=0.5}$    b1 0.585801 0.587922 0.590115 0.592865 0.594531
        a2 0.685801 0.686322 0.689115 0.691865 0.693431
        b2 0.296975 0.298123 0.300002 0.300925 0.301407
        $\bm{\zeta}$ 4.603475 4.604672 4.606924 4.608952 4.609963
        a1 0.956227 0.958796 0.959321 0.961059 0.962998
$\bm{\gamma=0.8}$    b1 0.588284 0.589483 0.591731 0.593654 0.595411
        a2 0.687284 0.689483 0.691731 0.693654 0.694411
        b2 0.294589 0.296798 0.298005 0.300588 0.302278
Further, the optimal behavior of the tuning parameter $\gamma$ is studied for pure data and contaminated data. In addition to finding the value of $\gamma$, the MSE of the model parameters is calculated under varying contamination levels with a fixed sample size of $m=100$. This same analysis has been conducted for both the GSM and IWJ methods for comparison purposes. Table 5 and Table 6 show the optimal tuning parameters identified through the GSM and IWJ approaches at $\bm{\epsilon}=0.000,0.006,0.044,0.085,0.153$. According to these optimal values of $\gamma$, the MSE of the model parameters is calculated to assess the performance of each method. The results reveal that the MSE obtained through the GSM method is significantly lower than that achieved using the IWJ method, indicating the superior accuracy and effectiveness of the GSM approach in identifying the optimal tuning parameter. Consequently, it can be inferred that GSM provides a more effective and reliable approach for finding the optimal tuning parameter $\gamma$ compared to the IWJ method.
Table 2: Estimated values of the parameters under Restricted case
        $\bm{\zeta}$ 4.597892 4.635432 4.693444 4.695126 4.719544
        a1 0.909448 0.916151 0.920746 0.926234 0.932904
MLE     b1 0.520801 0.529136 0.538825 0.550985 0.562547
        a2 0.644396 0.653997 0.664375 0.673456 0.681883
        b2 0.232281 0.242296 0.250991 0.259443 0.266687
        $\bm{\zeta}$ 4.600052 4.600212 4.600440 4.600891 4.601045
        a1 0.910248 0.911457 0.912640 0.913237 0.914001
$\bm{\gamma=0.2}$    b1 0.522408 0.523640 0.524965 0.525582 0.526946
        a2 0.645081 0.646897 0.647564 0.648336 0.649924
        b2 0.234308 0.236996 0.237980 0.239287 0.240867
        $\bm{\zeta}$ 4.600012 4.600000 4.600560 4.600011 4.600069
        a1 0.910528 0.911572 0.911999 0.913014 0.914872
$\bm{\gamma=0.5}$    b1 0.521265 0.522847 0.523921 0.524518 0.525889
        a2 0.646988 0.647671 0.648025 0.648253 0.649140
        b2 0.235658 0.236296 0.237274 0.238985 0.239423
        $\bm{\zeta}$ 4.600000 4.600156 4.600063 4.600215 4.600658
        a1 0.911289 0.912253 0.912945 0.913361 0.914125
$\bm{\gamma=0.8}$    b1 0.520836 0.521642 0.524275 0.525943 0.526596
        a2 0.646186 0.648639 0.649015 0.649541 0.650154
        b2 0.234975 0.235197 0.237625 0.238032 0.239413
6.2. Real Data Analysis
This research examines survey data from the National Health and Nutrition Examination Survey (NHANES) carried out from 2017 until March 2020. As a result of the COVID-19 pandemic, field operations for the 2019–2020 cycle were suspended, resulting in data that is incomplete and not representative of the nation. To tackle this issue, a nationally representative dataset was formed by merging data from the 2017–2018 cycle with pre-pandemic information from the 2019–March 2020 cycle. This data is retrieved from the adjacent link, National Health and Nutrition Examination Survey (NHANES); NHANES 2017-March 2020 Pre-pandemic Questionnaire Data - https://www.cdc.gov/nchs/nhanes. Here, the study centers on 50 subjects with equivalent follow-up durations, exploring $N_{i1l}$ and $N_{i2l}$, which represent the number of cold/flu and pneumonia events, correspondingly, in the period $(\tau_{l-1},\tau_{l}]$. These subjects have been observed every 3 months, over an entire period of 12 months.
@@ 6.2. Real Data Analysis
Table 3: MSE of estimates of parameters under Unrestricted case
        $\bm{\zeta}$ 0.009983 0.021282 0.039132 0.050652 0.061431
        a1 0.002659 0.004533 0.005967 0.007151 0.008087
MLE     b1 0.004597 0.005888 0.006484 0.007597 0.008521
        a2 0.004021 0.005782 0.006487 0.008123 0.009353
        b2 0.004233 0.005999 0.006650 0.007248 0.008747
        $\bm{\zeta}$ 0.010010 0.010112 0.012147 0.013074 0.015075
        a1 0.003218 0.003517 0.003742 0.004047 0.004513
$\bm{\gamma=0.2}$b1 0.004864 0.004981 0.005142 0.005535 0.005691
        a2 0.004266 0.005071 0.005256 0.005397 0.005788
        b2 0.004686 0.004781 0.005042 0.005651 0.006040
        $\bm{\zeta}$ 0.010045 0.010127 0.011147 0.012002 0.013142
        a1 0.003619 0.004218 0.004315 0.005017 0.005318
$\bm{\gamma=0.5}$b1 0.004081 0.004356 0.004605 0.004864 0.005251
        a2 0.004266 0.004923 0.005063 0.005071 0.005366
        b2 0.004389 0.004525 0.004887 0.005686 0.005991
        $\bm{\zeta}$ 0.010000 0.011015 0.012002 0.012089 0.013012
        a1 0.003619 0.004018 0.004123 0.004218 0.004417
$\bm{\gamma=0.8}$b1 0.004081 0.008864 0.009043 0.008864 0.008864
        a2 0.004661 0.004961 0.005134 0.005266 0.005371
        b2 0.004922 0.005312 0.005563 0.005686 0.005716
@@
        $\bm{\zeta}$ 0.009064 0.012095 0.025129 0.035234 0.038696
        a1 0.002824 0.004354 0.005814 0.006404 0.007299
MLE     b1 0.002223 0.003026 0.003538 0.004674 0.005206
        a2 0.002349 0.003304 0.004723 0.005711 0.006564
        b2 0.002691 0.004415 0.005463 0.006706 0.007087
        $\bm{\zeta}$ 0.010001 0.010097 0.011001 0.012142 0.013956
        a1 0.002948 0.002993 0.003220 0.003303 0.003802
$\bm{\gamma=0.2}$b1 0.002446 0.002783 0.003115 0.003656 0.004069
        a2 0.002637 0.002813 0.003161 0.003567 0.003656
        b2 0.003654 0.003799 0.004158 0.004399 0.004668
        $\bm{\zeta}$ 0.009856 0.010000 0.010161 0.010002 0.010443
        a1 0.003220 0.003419 0.003604 0.003918 0.004159
$\bm{\gamma=0.5}$b1 0.003115 0.003598 0.003897 0.004081 0.004266
        a2 0.003661 0.003768 0.003891 0.004166 0.004260
        b2 0.003358 0.003492 0.003716 0.003922 0.004169
        $\bm{\zeta}$ 0.009975 0.010000 0.010981 0.010000 0.010012
        a1 0.002619 0.002820 0.003012 0.003218 0.003433
$\bm{\gamma=0.8}$b1 0.002932 0.003198 0.003303 0.003598 0.003781
        a2 0.002673 0.002856 0.003017 0.003166 0.003366
        b2 0.002830 0.003058 0.003228 0.003458 0.003622
@@
        $\bm{\gamma}$ 0.58 0.34 0.41 0.47
        $\bm{\zeta}$ 0.000591 0.001717 0.000662 0.001447
        a1 0.001115 0.003672 0.001458 0.002965
MSEb1 0.000060 0.000908 0.000805 0.001261
        a2 0.000443 0.001190 0.000998 0.001330
        b2 0.000766 0.001393 0.000967 0.001857
The real dataset reveals a few outliers, particularly in the relationship between frequent cold/flu and the likelihood of developing pneumonia. In most cases, individuals who have experienced cold or flu episodes four to five times within a specific interval tend to develop pneumonia at least once in that period. However, 6 out of 50 individuals deviate from this pattern. Despite suffering from cold/flu five to six times in a given timeframe, they did not develop pneumonia at all. A key observation is that these individuals engage in a higher-than-average duration of physical activity, as depicted in the stacked bar graph (Figure 3), which visualizes the recurrence of cold/flu ($N_{i1l}$) and pneumonia ($N_{i2l}$) across individuals. Moreover, Figure 3 clearly illustrates that cold/flu episodes, represented by $N_{i1l}$, occur more frequently than pneumonia episodes, denoted by $N_{i2l}$. This observed pattern justifies the inequality constraints $a_{1}\geq a_{2}$ and $b_{1}\geq b_{2}$, which are imposed to reflect the higher recurrence rate of cold/flu compared to pneumonia in the model. The frailty variable $Z_{i}$, representing the number of hours a person engages in physical activities weekly, emerges as a crucial determinant in this anomaly. It appears that higher physical activity levels serve as a protective factor, preventing pneumonia during the follow-up period. This correlation is further illustrated in the bubble chart (Figure 4), where these six individuals are represented by large bubbles, reflecting their high exercise rates. Keeping in view the current scenario, the robust restricted minimum density power divergence (RMDPD) may be a suitable estimation technique.
Further, a dissimilarity measure-based test statistic to evaluate the suitability of the nonhomogeneous Poisson process for the data is defined as
$T=\frac{1}{mk}\sum_{i=1}^{m}\sum_{l=1}^{k}\left[\left|\frac{n_{i1l}-e_{i1l}}{e_{i1l}}\right|+\left|\frac{n_{i2l}-e_{i2l}}{e_{i2l}}\right|\right].$ (24)
@@
where $n_{i1l}$ and $e_{i1l}$ represent the observed and expected counts of cold or flu occurrences for the $i^{\text{th}}$ individual during the interval $(\tau_{l-1},\tau_{l}]$, respectively, and $n_{i2l}$ and $e_{i2l}$ represent the observed and expected counts of pneumonia occurrences for the same individual within the interval $(\tau_{l-1},\tau_{l}]$, with $i=1,2,3,\ldots,50$. The bootstrap (Bt) technique is employed to estimate the p-value of the test using 10,000 generated samples. A test statistic is computed for each sample. The proportion of bootstrap test statistics that exceed the test statistic from the real data, T, divided by the total number of bootstrap samples, provides the approximate p-value. A smaller test statistic indicates a greater likelihood that the data conform to the nonhomogeneous Poisson process. The computed values of the test statistic, T along with the associated p-values for various tuning parameters $(\gamma=0.2,0.5,0.8)$ are reported in Table 7. These p-values confirm the suitability of the nonhomogeneous Poisson process for modeling data. Therefore, it can be inferred that the instances of cold or flu and pneumonia in people can be adequately represented by the nonhomogeneous Poisson process.
@@
0.2 2.052768 0.510800
0.5 1.933244 0.654200
0.8 2.126994 0.359600
The MLE and RMDPDE of the model parameters are obtained, and the bootstrap estimates of Bias and MSE generated from 10,000 bootstrap samples are computed. The initial parameter value of the Sequential convex programming algorithm is set as (6.75, 0.82, 0.63, 0.42, 0.23), which is selected through a comprehensive grid search method. It is noticed that MDPDE shows lower Bootstrap Bias and MSE (Bt Bias and Bt MSE) in comparison to MLE.
Considering that the dataset contains some outliers, the efficacy of RMDPDE in this context aligns with the aim of the current study. Consequently, the restricted MDPDE has demonstrated outstanding performance on the real dataset. Table 8 displays the parameter estimates under both unrestricted and restricted settings. These results indicate that the inequality constraints, namely $a_{1}\geq a_{2}$ and $b_{1}\geq b_{2}$, are satisfied. The Bootstrap Bias and MSE (Bt Bias and Bt MSE) of MLE and RMDPDE for different $\gamma$ values are presented in Table 9 for the unrestricted case and in Table 10 for the restricted case. From these tables, it is evident that both MLE and RMDPDE exhibit a notable reduction in Bootstrap Bias and MSE under the restricted framework, highlighting the advantage of incorporating inequality constraints. This suggests that the restricted MDPDE offers superior performance among the estimators considered. The optimal tuning parameter, $\gamma$, for the model, is determined using real data. Based on this value of $\gamma$, the model parameters $\zeta$, $a_{1}$, $b_{1}$, $a_{2}$, and $b_{2}$ are estimated, and the bootstrap bias and mean squared error (MSE) of the parameters are calculated. The results obtained using GSM method are compared with those from the IWJ method. Table 11 provides a detailed summary of the optimal tuning parameter $\gamma$, the estimated model parameters, and their corresponding Bt Bias and Bt MSE values. It has been noted that the estimates derived from the GSM method consistently show lower Bt Bias and Bt MSE compared to those obtained from the IWJ method. This comparison reveals that the GSM method outperforms the IWJ method by achieving lower MSE, underscoring its effectiveness in tuning parameter selection.
Table 8: Estimated values of the parameters under Unrestricted and Restricted cases
@@
Unrestricted MLE 6.872250 0.914841 0.725169 0.330802 0.193911
$\bm{\gamma=0.2}$ 6.766742 0.844871 0.655538 0.406642 0.218145
$\bm{\gamma=0.5}$ 6.761689 0.840328 0.652417 0.404723 0.213945
$\bm{\gamma=0.8}$ 6.739611 0.803824 0.613709 0.409200 0.219717
Restricted MLE 6.799012 0.829115 0.638724 0.425638 0.232427
$\bm{\gamma=0.2}$ 6.752852 0.822441 0.632660 0.423361 0.231178
$\bm{\gamma=0.5}$ 6.752557 0.821889 0.632196 0.422114 0.231008
$\bm{\gamma=0.8}$ 6.751632 0.821745 0.631684 0.421248 0.230181
Table 9: Bootstrap Bias and MSE of estimates of parameters under Unrestricted case
$\bm{\zeta}$ a1 b1 a2 b2
Bt Bias MLE 0.122140 0.094960 0.095150 -0.089300 -0.036238
$\bm{\gamma=0.2}$ 0.016694 0.024955 0.025646 -0.013667 -0.011995
$\bm{\gamma=0.5}$ 0.011575 0.020198 0.022350 -0.015346 -0.016190
$\bm{\gamma=0.8}$ -0.010576 -0.016552 -0.016488 -0.010609 -0.010283
Bt MSE MLE 0.049346 0.009017 0.008864 0.005825 0.002493
$\bm{\gamma=0.2}$ 0.002993 0.002345 0.002416 0.003212 0.001096
$\bm{\gamma=0.5}$ 0.002491 0.001907 0.002012 0.002017 0.001008
$\bm{\gamma=0.8}$ 0.001715 0.001808 0.001894 0.001591 0.000945
Table 10: Bootstrap Bias and MSE of estimates of parameters under Restricted case
$\bm{\zeta}$ a1 b1 a2 b2
Bt Bias MLE -0.024464 -0.035629 -0.021732 -0.040737 -0.009072
$\bm{\gamma=0.2}$ 0.001589 0.002138 0.002899 -0.003837 -0.002089
$\bm{\gamma=0.5}$ 0.001289 0.001744 0.002390 -0.003081 -0.001932
$\bm{\gamma=0.8}$ -0.001943 -0.001389 -0.001938 -0.003128 -0.001867
Bt MSE MLE 0.002436 0.003273 0.004123 0.002287 0.000931
$\bm{\gamma=0.2}$ 0.000199 0.000136 0.000243 0.000172 0.000087
$\bm{\gamma=0.5}$ 0.000161 0.000112 0.000204 0.000141 0.000085
$\bm{\gamma=0.8}$ 0.000163 0.000148 0.000433 0.000158 0.000035